OpenStudy (qaisersatti):

FINALLY COMPLETED THE END TERM...WITH ALL YOUR HELP I WAS ABLE TO SECURE 100% RESULT THANKS GUYS...I WILL WELCOME ANY QUESTIONS IF ANYONE HAS....THANKS

OpenStudy (nurali):

Congratulation.

OpenStudy (anonymous):

plz help at Q6

OpenStudy (anonymous):

List out the formulas with right answers for your values will u? this will help a lot of us in a more verified way. Thnx

OpenStudy (qaisersatti):

thats gonna take time actually...if i write the answers :) but i will do it tonight @ nurali thanks brother.

OpenStudy (anonymous):

this will help us brother

OpenStudy (anonymous):

Congrats :-)

OpenStudy (anonymous):

Please note that they often change sequestioon of parts in questions

OpenStudy (qaisersatti):

Q1 a)ia=2*i1*R/(2*R+rm) and vo=ia*rm b)iA =(-i2)*R/(2*R+rm) and V0=ia*rm c) part a) + part b) d)2R Q2 p1: Vs as vin<vt p2: v0 = vs-kr(e^((vin-vt)/vn)-1) p3: vin = ln(vs/2kr +1)vn+vt p4: vin=sqrt(vs/kr)+vt p5: -kre^((vin-vt)/vn) p6: -kr(vin-vt) Q 3 SAME AS HOMEWOR 8 The answers are as follows if you don’t want to solve it  0 2000 270.6 3030.3 146.36 Q4 p1=V*R2/(R1+R2)*(1-e^(-(R1+R2)*t/(R1*R2*C))) p2=(V*R2/(R1+R2))*cos(1/sqrt(L*C)*(t-T)) p3=C*(V*R2/(R1+R2))*sin((t-T)*1/sqrt(L*C))/sqrt(L*C) Q5 E V B Y F Z Q6 A) L/(Rw+RL) B) L/(Rw+RL) C) RL×V0×(1−e^−T×(Rw+RL)/L)(Rw+RL)×11−(e^−T×(Rw+RL)/L)2 D) RL×V0×(1−e^−T×(Rw+RL)/L)(Rw+RL)×11−(e^−T×(Rw+RL)/L)^2×e−T×(Rw+RL)/L Good luck 

OpenStudy (anonymous):

@Qaisersatti, plzzz help! Q6, my values are... V0=5V, T=0.1ms, Rw=1Ω, Lw=0.5mH and RL=4Ω.